my qusestion issssssssssssssssssssssssssssssssssssssssssssssss
 
 
 
if circle x2+y2-kx-12y+4=0 touches x axis then k=
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 

Grade:12th pass

1 Answers

poonam
6 years ago
General equation of circle is x2+y2+2gx+2fy+c=0. So comparing this equation with the equation in the question we get g=-k/2. f=-6. C=4. As the circle touches x axis so G^2=c. So k^2/4=4. k=4